OpenStudy (anonymous):

Let P(x) = 6x − 7. Find the value. P(x + h) − P(x)

OpenStudy (anonymous):

6h

OpenStudy (anonymous):

P(x+h)=6(x+h)-7=6x+6h-7 P(x)=6x-7 P(x+h)-P(x)=(6x+6h-7)-(6x-7) =6x+6h-7-6x+7 =6h